OpenStudy (anonymous):

Can somebody please help me. |n/6|+3 = 4 I don't understand this?

OpenStudy (timmyk14):

Distributive property I believe

OpenStudy (anonymous):

Its absolute value. my teacher doesn't really explain things

OpenStudy (johnweldon1993):

First things first...you want that absolute value alone on one side of the equation...subtract 3 from both sides...

OpenStudy (anonymous):

subtract the 3 from the 4?

OpenStudy (timmyk14):

3*n...3*6 3n+18=4 3n=-14 n~-4.7

OpenStudy (timmyk14):

I think

OpenStudy (johnweldon1993):

Right... \[\large |\frac{n}{6}| + 3 = 4\] \[\large |\frac{n}{6} | = 1\] Now...you know that in an absolute value ...you can never get a negative number....so this an be re-written as \[\large \frac{n}{6} = 1\] or \[\large \frac{n}{6} = -1\] There will be 2 solutions here...

OpenStudy (anonymous):

ohhh okay I get it. so 6 or -6?

OpenStudy (johnweldon1993):

Correct @glitterythings

OpenStudy (anonymous):

thanks :)

OpenStudy (timmyk14):

Oops didn't see the absolute value signs